Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR21 @ 14:00 pm on January 24th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • CC Magnet status
  • Hydrogen status
  • Issues arising from previous running
  • Shutdown work in Jan and Feb: present version
  • Operations :
    • C&M System Check status (Ajit/Ed)
    • Establishing a timetable for the runup : when should systems (including magnets) be ready for running?
      • Beam schedule
        • Run up on Jan 30 - Feb 12.
        • Start of user cycle : Feb 14
        • Start of data-taking : Feb 15
      • Reserve shifter support from Feb 10 - Feb 14 (Includes a weekend though)
      • MOM formally starts on Feb 9 (but should be present for the runup activities but will not be on 24-hour call)
      • Possible runup plan (framework for discussion)
        • Bring up magnets to running configuration, take some pulser data and take magnets down (1 day during week of Feb 6 - 9).
          • Validate and lock C&M.
        • Survey and test detectors in the week of Feb 6 - 9
        • Expert led run on Fri Feb 10 (with/without magnets (?))
        • Triage actions of expert led run on Mon Feb 13
        • Bring up magnets on Feb 14, with overnight monitoring by reserve shift.
        • Data-taking start at 08:00 on Feb 15
  • Shift signup for first half of Feb is viable now (but not full)
Shift Block ID # signed up # shifters still needed
1 6 0
2 4 2
3 4 2
4 4 2
Note that in the last 3 shift blocks we still need at least one more shifter overnight for magnet
monitoring. Otherwise we will have to rearrange the shift and take only 1 data-taking shift with 2
magnet monitoring shifts.
  • BLOCs are needed
  • AOB:

Minutes

  • CC Magnet status (JBoehm)
    All 3 CC magnets are now in flip mode; cold and full of helium. Everything is now being run into shorts for the CC IOC testing.
    Doing dry runs into shorts to learn how to run SS magnets from "Central CC GUI"; writing expert-level instructions.
    "Central CC GUI" is expert-level, to be run by experts only. There was some discussion as to whether it should be restricted to RR2 (to avoid accidental shifter interactions) or run from the MLCR where much more information is available. Decide in Controls Review.
    Will soon need to run magnets for real: after next week when Daresbury are down.
  • Hydrogen status (JBoehm)
    Got agreement from Steve Harrison and have restarted the R9 LH2 system today with Neon: Mark Tucker starting it now.
    Cooling slowly as the cooling is pressure dependent.
    Have cover until the end of the week: plan depends on cooling rate.
  • Issues arising from previous running (DC)
    Need a comprehensive list of encountered issues, else can't assume they will be addressed before running resumes. We are now well past halfway through the shutdown.
  • Shutdown work in Jan and Feb ( official version )
    Issues/Progress on workplan items:
    • Asymmetric ramping of FC - ramp-down is faster: there are 3 stages on the way up but only 2 on way down; caused a mild pressure build-up. Being investigated why PSU behaved differently on the way up and way down.
    • Paul Hodgson has checked with MOMs - no further issues have been identified.
    • POMPOMs sensor test - imminent
    • Proton Absorbers - in progress
    • "New" chiller plumbing - done
    • #1878 Trench heat exchanger to counter-flow - done
    • Diffuser iris 4: needs next SSU run, week of 6th. Too late to remove mechanism now anyway: only option if the problem remains is to lock that iris open (or closed) for the duration of the User Cycle
    • #1399 MICEnet lockout - tomorrow morning
    • #1888 Diffuser air - reported done
    • [[cm:SystemSpreadsheets|State Machines]]:
      • #1896 Target - in progress
      • #1880 Beamline - in progress
      • #1881 Services and Environment - in progress. PVs to be "clones" to allow system-specific limits to be applied via the systems' own spreadsheet. Services and Environment sheet to catch the rest. Compressed air PV needs cloning for the Decay Solenoid. MICE Hall temperature should be limited to 35°C to avoid problems with control systems. No feedback yet from overwhelming majority of subsystems.
      • #1881 Computing - includes UPSs #1213 and File Compactor #1895. Has been given to Paolo.
    • Macro "Code Topology" - in progress
    • Magnet overview for shifters - ???
    • Electrical: DL down en masse week of 30th
    • NIKHEF Hall Probes
      There are 4 on each SS. 3 on SSU and 2 on SSD are OK, other 3 are duff - failed as switched on/off incorrectly.
      They are needed by the Tracker to figure out particle momentum; Channel IOC depends on them when calculating current and configuration (solenoid vs flip); and DAQ needs PV as it writes them in the data-stream. This all breaks every time another probe is lost.
      Proposed to remove the Hall probes from the Cooling Channel IOC and write a new, dedicated IOC. Ajit to plan this.
      In the meantime remove the dependencies on the presently-failed one.
    • Presently still a lot of ALH activations going off
  • Shutdown work in Jan and Feb:
    The PPS functional test will be on Wed. 25th January - MICE Hall closed.
    Controls Testing Thu/Fri 25th/26th
    ODH sensor service will be on 30th January.
    DL down en masse week 30th January for electrical work.
    Diffuser test needs to be re-scheduled for February
    CC Magnet tests week of 6th
    Ramp CC magnets morning of 14th and watch them overnight (soak test); start data-taking next morning (15th).

Actions

  • JBoehm to get SS operating info from PHanlet (ongoing)
  • JBoehm to get/impose consensus from SC Magnets group on running of "Central CC GUI"
  • DC to collate list of issues arising from previous running
  • Status of miceraid5 disk cleanup (stands)
  • AKurup to make and circulate plan for NIKHEF probe IOC
  • AKurup to make and circulate plan for C&M cold-start test
  • SBoyd - ISIS machine-physics beam time
  • DC to ensure Health Physics have checked temporary neutron monitor (stands)
  • SBoyd - ensure MICE-wide discussion of MOM C&M panel

Updated by Nebrensky, Henry over 6 years ago · 21 revisions